# A Bayesian Estimator of Sample Size

### v2

@@ Abstract
We consider a Bayesian estimator of sample size (BESS) and an application to oncology dose optimization clinical trials. BESS is built upon three pillars, Sample size, Evidence from observed data, and Confidence in posterior inference. It uses a simple logic of “given the evidence from data, a specific sample size can achieve a degree of confidence in the posterior inference.” The key distinction between BESS and standard sample size estimation (SSE) is that SSE, typically based on Frequentist inference, specifies the true parameters values in its calculation while BESS assumes possible outcome from the observed data. As a result, the calibration of the sample size is not based on type I or type II error rates, but on posterior probabilities. We demonstrate that BESS leads to a more interpretable statement for investigators, and can easily accommodates prior information as well as sample size re-estimation. We explore its performance in comparison to the standard SSE and demonstrate its usage through a case study of oncology optimization trial. BESS can be applied to general hypothesis tests. An R tool is available at https://ccte.uchicago.edu/BESS.
@@ 1.1 Motivation
We consider a Bayesian framework for SSE. The research is motivated by the recent change in early-phase oncology drug development that recommends randomized comparison of multiple doses. In the past two decades, oncology drug development has benefited from biological and genomics breakthroughs and novel cancer drugs are no longer based on cytotoxic one-size-fits-all mechanism like chemicals or radiations. Instead, targeted, immune, and gene and cell therapies combat tumor cells by precisely altering oncogenic cellular or molecular pathways. Consequently, the traditional monotonic dose-response relationship is no longer valid for many novel cancer drugs. Instead, efficacy often plateaus or even decreases after dose rises above certain level. To this end, US FDA launched Project Optimus (Food et al., 2023; Shah et al., 2021; Blumenthal et al., 2021) aiming to transform the early-phase oncology drug development. The main objective of the project is to identify an optimal dose, potentially lower than the maximum tolerated dose but with at least comparable anti-tumor effect. The new dose optimization paradigm recommends a randomized comparison of two or more doses following an initial dose escalation. While randomized trial design and SSE are routinely conducted in drug clinical trials, they are new in dose-optimization trials, usually restricted by limited resources. Type I/II error rates for standard SSE are rarely the main objective for early-phase of drug development. Instead, investigators are more concerned about the accuracy in the trial decision making, such as selecting the right dose or starting confirmatory studies, which may be very costly. Current SSE for dose optimization trials is often based on ad-hoc choices, such as using a small sample size of 20 patients per dose. This leads to a question of how much is expected to learn from the 20 patients, or whatever the number may be. Moreover, clinical trials that use human subjects for clinical research should always provide a reasonable justification for the number of subjects to be enrolled. We attempt to fill the gap by considering a Bayesian approach for sample size estimation.
@@ 1.2 Review of SSE Methods
In the literature, standard SSE methods (Adcock, 1997; Wittes, 2002; Desu, 2012) are based on Frequentist hypothesis testing by controlling the type I error rate $\alpha$ to achieve a desirable power $(1-\beta)$, assuming the true population parameters are known. For example, a sample size statement for a two-arm RCT based on binary outcome is as follows:
  • Statement 1: At type I error rate of $\alpha$, with a clinically minimum effect size $\theta^{*}$, $S$ subjects are needed to achieve $(1-\beta)$ power when the response rates for the treatment and control arms are $\theta_{1}$ and $\theta_{0}$.
A comprehensive review can be found in Wang and Ji (2020). In addition, sample size may be determined based on hybrid Frequentist-Bayesian inference as in Ciarleglio and Arendt (2017); Berry et al. (2010). These methods use Bayesian models but calibrate sample size based on type I error rate and power via simulation. In contrast, some methods use Bayesian properties for sample size estimation such as the length, coverage of posterior credible intervals, Bayes factors under sampling and fitting priors (Lin et al., 2022), or loss functions like $k\cdot\text{False Positive Rate}+\text{False Negative Rate}$ (Müller et al., 2004).
@@ 1.3 Main idea
To this end, we propose a new Bayesian estimator of sample size (BESS) based on a general hierarchical modeling framework and posterior inference. Let $\bm{y}$ denote the observed data and $\bm{\theta}$ the unknown parameters as generic notation. We assume a Bayesian hierarchical model is to be used for the experimental design and data analysis. Our main idea is originated from noticing a simple trend between the sample size and observed effect size from data. Suppose in a two-arm RCT with sample size $n$ per arm and a binary outcome, we are interested in testing if the effect size (difference in response rates), defined as $\theta=(\theta_{1}-\theta_{0})$ between the treatment (1) and control (0) is greater than a clinically minimum effect size $\theta^{*}$. This can be formulated as testing the null hypothesis $H_{0}:\theta\leq\theta^{*}$ versus the alternative $H_{1}:\theta>\theta^{*}.$ Suppose after the trial is completed, the observed response rates are $\bar{y}_{1}$ and $\bar{y}_{0}$ for the treatment and control arms, respectively. A Frequentist $z$-test is used to test the hypotheses given by
@@
Instead of using a Frequentist inference like the $z$-test, we consider Bayesian hypothesis testing based on posterior probability of the alternative hypothesis, which, in the setting of clinical trials, refers to the success of the trial. Decision of accepting the alternative hypothesis is by thresholding the posterior probability of $H_{1}$ at a relatively large value $c\in(0,1)$. Assume a Bayesian hierarchical model is given by $f(\bm{y}\mid\bm{\theta})\pi(\bm{\theta}\mid H)\pi(H)$, where $H=H_{0}$ or $H_{1}$ is a binary indicator of the null and alternative hypotheses. The proposed BESS aims to find a balance between 1) “Sample size” $S$ of the clinical trial, 2) “Evidence” defined as the observed treatment effect $e=(\bar{y}_{1}-\bar{y}_{0})$, and 3) “Confidence” defined as the posterior probability of the alternative hypothesis.
@@
  • Statement 2: Assuming the evidence is at least $e$, $S$ subjects are needed to declare with confidence $c$ that the treatment effect is at least $\theta^{*}$.
Note that “evidence” is a function of the data, not parameters, and “confidence” refers to $\text{Pr}(H_{1}|\bm{y})$, the posterior probability of the alternative hypothesis, i.e., the treatment effect is at least $\theta^{*}.$ Comparing Statement 1 and Statement 2, one can see that the two statements are based on different statistical properties, Frequentist type I/II error rates for Statement 1 and Bayesian posterior probabilities for Statement 2. Also, while Statement 1 assumes the true values of the parameters are given, Statement 2 assumes what might be observed from the trial data. We will show that the BESS and associated Statement 2 is easier to interpret in practice since it directly addresses the uncertainty in the decision to be made for the trial at hand, measured by posterior probabilities.
@@ 2 Probability Model
We consider BESS for both one-arm and two-arm trials. Denote $y_{ij}$ the outcome of patient $i$ in arm $j$, where $i=1,\ldots,n,$ index the patients, and $j=0$ and $1$ index the control and treatment arm, respectively. When needed, we drop index $j$ and use $y_{i}$ for one-arm trials. Let $\theta_{1}$ denote the true response parameter for the treatment arm, and let $\theta_{0}$ be the true response parameter for the control arm in a two-arm trial or the reference response parameter in a one-arm trial. Let $\theta=f(\theta_{0},\theta_{1})$ a function of $\theta_{0}$ and $\theta_{1}$ denote the treatment effect. For example, we consider $\theta=(\theta_{1}-\theta_{0})$ in this paper although its form can be generalized. Consider hypotheses
@@
$y_{ij}|\theta_{j}$ $\sim$ $f(\theta_{j}),$
$(\theta_{0},\theta_{1})|\tilde{\bm{\theta}},H=H_{j}$ $\sim$ $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})I(\theta\in H_{j}),$ (2)
@@
where $f(.)$ represents the likelihood function, $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})$ is a joint prior distribution for $(\theta_{0},\theta_{1})$, $\tilde{\bm{\theta}}$ are hyper-parameters, and $I(x\in A)$ is the indicator function which equals 1 if $x\in A$ and 0 otherwise. For simplicity, we consider $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})=\pi_{0}(\theta_{0}|\tilde{\bm{\theta}})\pi_{1}(\theta_{0}|\tilde{\bm{\theta}})$, where $\pi_{0}=\pi_{1}$.
In this work, we consider three specific types of outcome $y_{ij}$: binary, continuous, and count. A summary of the parameter $\theta_{j}$, likelihood function $f(.)$, and prior distribution $\pi_{j}$ for $\theta_{j}$ is shown in Table 1. We consider conjugate prior for simplicity and computational convenience although the proposed BESS works for any general priors. While it is not the focus of this work to discuss choice of priors, we note the flexibility of BESS to incorporate various priors in real-life applications. For example, when little prior information is known vague priors like Jeffrey’s prior may be considered; in contrast, it is also possible to use informative priors when prior information is available. For instance, assume a previous trial with binary outcome has completed with a sample size of $n_{0}$ patients of whom their outcome data are available, denoted as $\bm{y}^{0}=\{y_{i}^{0};i=1,\ldots,n_{0}\}$. Then, one could consider an informative prior $\pi_{1}(\theta_{1})=\text{Beta}(a,b)$ and set $a=a^{*}+\sum_{i=1}^{n_{0}}y_{i}^{0},\,\,b=b^{*}+n_{0}-\sum_{i=1}^{n_{0}}y_{i}^{0},$ where $a^{*}$ and $b^{*}$ are small (e.g., $a^{*}=b^{*}=0.5$).
Outcome type Parameter $\theta_{j}$ Likelihood $f(\cdot)$ Prior distribution $\pi_{j}(\tilde{\bm{\theta}})$
@@ 3.1 Confidence
We introduce the three pillars of BESS, Confidence, Evidence, and Sample Size. We start with “Confidence”, the confidence of posterior inference, expressed mathematically as the posterior probability of the alternative hypothesis $\text{Pr}(H=H_{1}|\bm{y}_{n}),$ where $\bm{y}_{n}$ denotes the data with sample size $n$. The optimal decision rule under a variety of loss functions (Müller et al., 2004) is to reject the null $H_{0}$ and accept $H_{1}$ if $\text{Pr}(H=H_{1}|\bm{y}_{n})\geq c$ for a high value of $c\in(0,1).$ The higher $c$ is, the more confident is the decision. To see this, one simply observes that $(1-c)$ is the upper bound of posterior probability of a wrong rejection since when $H_{0}$ is rejected, $\text{Pr}(H=H_{0}|\bm{y}_{n})<c$.
@@ 3.2 Evidence
In simple words, evidence $e$ is the observed effect size from the trial data before they are observed. This means that in order to apply BESS, investigator needs to pre-specify (and calibrate) the potential observed effect size before the trial is conducted. This is analogous to the requirement of specifying the true parameter values in the standard SSE, except BESS assumes what might be observed rather than what might be true. We consider evidence $e$ as a function of the sufficient statistic in Table 1. For the three outcomes in one-arm trials, $(e+\theta_{0})$ is exactly the sufficient statistic. For two-arm trials, $e$ is the difference between the sufficient statistics $(\bar{y}_{0},\bar{y}_{1})$ of the two arms. Being a function of sufficient statistics allows for a search algorithm to find an appropriate sample size under BESS, which will be clear next.
@@ 3.3 Sample Size of BESS
For binary and count data in two-arm trials, $\text{Pr}(H=H_{1}|\bm{y}_{n})=\text{Pr}(H=H_{1}|\bar{y}_{1},\bar{y}_{0},n)$. See Appendix A.3 for a quick proof. Since $e=(\bar{y}_{1}-\bar{y}_{0})$, fixing evidence $e$ does not uniquely define $\text{Pr}(H=H_{1}|\bm{y}_{n}).$ Hence, we propose to specify evidence $e$ and find all pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$ that satisfies $e=(\bar{y}_{1}-\bar{y}_{0})$. We then find the minimum posterior probability of $H_{1}$ among these pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$, i.e.,
@@
Then the sample size statement for a two-arm trial is the same as BESS 1. And BESS Algorithm 2 provides details of finding the sample size $n$ for a two-arm trial.
BESS Algorithm 2 Two-Arm Trials; Binary or Count Data
@@
Alternatively, one may wish to specify the values of $\bar{y}_{1}$ and $\bar{y}_{0}$ directly instead of their difference $e$. Since $(\bar{y}_{1},\bar{y}_{0})$ is sufficient, the sample size search algorithm becomes easier. To this end, we propose a simple variation BESS Algorithm 2’ in Appendix A.4 and statement BESS 2.2 assuming $(\bar{y}_{1},\bar{y}_{0})$ are given.
@@
Lastly, since binary and count data are integer-valued, not all specified evidence value $e$ can be achieved for a given sample size in the proposed BESS Algorithms 1 and 2. For example, when the outcome is binary and sample size $n=10$, it is impossible to observe evidence $e=0.15$ since this requires having $n\cdot e=1.5$ more responders in the treatment arm than the control. The number of responders cannot be a fraction. To this end, we propose to round down the specified $e$ to the nearest possible value for a given $n$ by $e^{\prime}=\frac{1}{n}\lfloor n\cdot e\rfloor,$ where $\lfloor$.$\rfloor$ . ⌋ is the floor function. This gives a conservative estimate of the sample size since even with smaller evidence, the sample size would still ensure the needed confidence.

### v3

@@ Abstract
We consider a Bayesian estimator of sample size (BESS) and an application to oncology dose optimization clinical trials. BESS is built upon three pillars, Sample size of the experiment, Evidence from the observed data, and Confidence in the posterior inference. It uses a simple logic of “given the evidence from data, a specific sample size can achieve a degree of confidence in the posterior inference.” The key distinction between BESS and standard sample size estimation (SSE) is that SSE, typically based on Frequentist inference, specifies the true parameters values in its calculation while BESS assumes possible outcome from the observed data. As a result, the calibration of the sample size is not based on type I or type II error rates, but on posterior probabilities. We demonstrate that BESS leads to a more interpretable statement for investigators, and can easily accommodates prior information as well as sample size re-estimation. We explore its performance in comparison to the standard SSE and demonstrate its usage through a case study of oncology optimization trial. An R tool is available at https://ccte.uchicago.edu/BESS.
@@ 1.1 Motivation
The proposed research is motivated by the recent change in early-phase oncology drug development that recommends randomized comparison of multiple doses. In the past two decades, oncology drug development has benefited from biological and genomics breakthroughs and novel cancer drugs are no longer based on cytotoxic one-size-fits-all mechanism like chemicals or radiations. Instead, targeted, immune, and gene and cell therapies combat tumor cells by precisely altering oncogenic cellular or molecular pathways. Consequently, the traditional monotonic dose-response relationship is no longer valid for many novel cancer drugs. Instead, efficacy often plateaus or even decreases after dose rises above certain level. To this end, US FDA launched Project Optimus (USFDA, 2023; Shah et al., 2021; Blumenthal et al., 2021) aiming to transform the early-phase oncology drug development. The main objective of the project is to identify an optimal dose, potentially lower than the maximum tolerated dose but with at least comparable anti-tumor effect. The new dose optimization paradigm recommends a randomized comparison of two or more doses following an initial dose escalation. While randomized trial design and SSE are routinely conducted in drug clinical trials, they are new in dose-optimization trials, usually restricted by limited resources. Type I/II error rates for standard SSE are rarely the main objective for early-phase of drug development. Instead, investigators are more concerned about the accuracy in the final decision including selection of the right dose or launching future confirmatory studies, both of which may be very costly. Current SSE for dose optimization trials is often based on ad-hoc choices, such as using a small sample size of 20 patients per dose. This leads to a question of how much is expected to learn from the 20 patients, or whatever the number may be. Moreover, clinical trials that use human subjects for clinical research should always provide a reasonable justification for the number of subjects to be enrolled. We attempt to fill the gap by considering a Bayesian approach for sample size estimation.
@@ 1.2 Review of SSE Methods
In the literature, standard SSE methods (Adcock, 1997; Wittes, 2002; Desu, 2012) are based on Frequentist hypothesis testing aiming to control the type I error rate $\alpha$ and achieve a desirable power $(1-\beta)$, assuming the true population parameters are known. For example, a sample size statement for a two-arm RCT based on binary outcome is as follows:
  • Statement 1: At type I error rate of $\alpha$, with a clinically minimum effect size $\theta^{*}$, $n$ subjects are needed to achieve $(1-\beta)$ power when the response rates for the treatment and control arms are $\theta_{1}$ and $\theta_{0}$.
A comprehensive review can be found in Wang and Ji (2020). In addition, sample size may be determined based on hybrid Frequentist-Bayesian inference (Ciarleglio and Arendt, 2017; Berry et al., 2010). These methods use Bayesian models but calibrate sample size based on type I error rate and power via simulation. In contrast, some methods use Bayesian properties for sample size estimation such as the length, coverage of posterior credible intervals, Bayes factors under sampling and fitting priors (Chen et al., 2011; Lin et al., 2022; Chen et al., 2023), or loss functions like $k\cdot\text{False Positive Rate}+\text{False Negative Rate}$ (Müller et al., 2004). Novel Bayesian SSE approaches have also considered borrowing information from historical data (Chen et al., 2011; Yuan et al., 2022), mostly targeting device clinical trials.
@@ 1.3 Main idea
To this end, we propose a new Bayesian estimator of sample size (BESS) based on a general hierarchical modeling framework and posterior inference. Motivated by but aiming to simplify the work in Chen et al. (2011), we consider a more practical and simpler framework. Let $\bm{y}$ denote the observed data and $\bm{\theta}$ the unknown parameters as generic notation. We assume a Bayesian hierarchical model is to be used for the experimental design and data analysis. Our main idea is originated from noticing a simple trend between the sample size and observed effect size from data. Suppose in a two-arm RCT with sample size $n$ per arm and a binary outcome, we are interested in testing if the effect size (difference in response rates), defined as $\theta=(\theta_{1}-\theta_{0})$ between the treatment (1) and control (0) is greater than a clinically minimum effect size $\theta^{*}$. This can be formulated as testing the null hypothesis $H_{0}:\theta\leq\theta^{*}$ versus the alternative $H_{1}:\theta>\theta^{*}.$ Suppose after the trial is completed, the observed response rates are $\bar{y}_{1}$ and $\bar{y}_{0}$ for the treatment and control arms, respectively. A Frequentist $z$-test is used to test the hypotheses given by
@@
Instead of using a Frequentist inference like the $z$-test, we consider Bayesian hypothesis testing based on posterior probability of the alternative hypothesis, which, in the setting of clinical trials, refers to the success of the trial. Decision of accepting the alternative hypothesis is by thresholding the posterior probability of $H_{1}$ at a relatively large value $c\in(0,1)$. Assume a Bayesian hierarchical model is given by $f(\bm{y}\mid\bm{\theta})\pi(\bm{\theta}\mid H)\pi(H)$, where $H=H_{0}$ or $H_{1}$ is a binary indicator of the null and alternative hypotheses. The proposed BESS aims to find a balance between 1) “Sample size” $n$ of the clinical trial, 2) “Evidence” defined as the observed treatment effect $e=(\bar{y}_{1}-\bar{y}_{0})$, and 3) “Confidence” defined as the posterior probability of the alternative hypothesis.
@@
  • Statement 2: Assuming the evidence is at least $e$, $n$ subjects are needed to declare with confidence $c$ that the treatment effect is at least $\theta^{*}$.
Note that “evidence” is a function of the data, not parameters, and “confidence” refers to $\text{Pr}(H_{1}|\bm{y})$, the posterior probability of the alternative hypothesis, i.e., the treatment effect is at least $\theta^{*}.$ Comparing Statement 1 and Statement 2, one can see that the two statements are based on different statistical properties, Frequentist type I/II error rates for Statement 1 and Bayesian posterior probabilities for Statement 2. Also, while Statement 1 assumes the true values of the parameters are given, Statement 2 assumes what might be observed from the trial data. Alternatively, one may consider a predictive distribution of evidence by integrating the evidence over a prior of $\pi(\theta)$ as suggested in Kunzmann et al. (2021). However, the type of approaches would inevitably invoke the probability inference based on the sampling or marginal distribution of $\bm{y}$, deviating from our focus using Bayesian inference based on posterior distributions. We will show that the BESS and associated Statement 2 is easier to interpret in practice since it directly addresses the uncertainty in the decision to be made for the trial at hand, measured by posterior probabilities.
@@ 2 Probability Model
We consider BESS for both one-arm and two-arm trials. Denote $y_{ij}$ the outcome of patient $i$ in arm $j$, where $i=1,\ldots,n,$ index the patients, and $j=0$ and $1$ index the control and treatment arm, respectively. When needed, we drop index $j$ and use $y_{i}$ for one-arm trials. Let $\theta_{1}$ denote the true response parameter for the treatment arm, and let $\theta_{0}$ be the true response parameter for the control arm in a two-arm trial or the reference response parameter in a one-arm trial. Let $\theta=d(\theta_{0},\theta_{1})$ a function of $\theta_{0}$ and $\theta_{1}$ denote the treatment effect. For example, we consider $\theta=(\theta_{1}-\theta_{0})$ in this paper although its form can be generalized. Consider hypotheses
@@
$y_{ij}|\theta_{j}$ $\sim$ $l(\theta_{j}),$
$(\theta_{0},\theta_{1})|\tilde{\bm{\theta}},H=H_{j}$ $\sim$ $\frac{1}{C_{j}}\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})I(\theta\in H_{j}),$ (2)
@@
[…]
In this work, we consider three specific types of outcome $y_{ij}$: binary, continuous, and count. A summary of the parameter $\theta_{j}$, likelihood function $l(.)$, and prior distribution $\pi_{j}$ for $\theta_{j}$ is shown in Table 1. We consider conjugate prior for simplicity and computational convenience although the proposed BESS works for any general priors. While it is not the focus of this work to discuss choice of priors, we note the flexibility of BESS to incorporate various priors in real-life applications. For example, when little prior information is known vague priors like Jeffrey’s prior may be considered; in contrast, it is also possible to use informative priors when prior information is available. For instance, assume a previous trial with binary outcome has completed with a sample size of $n_{0}$ patients of whom their outcome data are available, denoted as $\bm{y}^{0}=\{y_{i}^{0};i=1,\ldots,n_{0}\}$. Then, one could consider an informative prior $\pi_{1}(\theta_{1})=\text{Beta}(a,b)$ and set $a=a^{*}+\sum_{i=1}^{n_{0}}y_{i}^{0},\,\,b=b^{*}+n_{0}-\sum_{i=1}^{n_{0}}y_{i}^{0},$ where $a^{*}$ and $b^{*}$ are small (e.g., $a^{*}=b^{*}=0.5$).
Outcome type Parameter $\theta_{j}$ Likelihood $l(\cdot)$ Prior distribution $\pi_{j}(\tilde{\bm{\theta}})$
@@ 3.1 Confidence
We introduce the three pillars of BESS, Confidence, Evidence, and Sample Size. We start with “Confidence”, the confidence of posterior inference, expressed mathematically as the posterior probability of the alternative hypothesis $\text{Pr}(H=H_{1}|\bm{y}_{n}),$ where $\bm{y}_{n}$ denotes the data with sample size $n$. The optimal decision rule under a variety of loss functions (Müller et al., 2004) is to reject the null $H_{0}$ and accept $H_{1}$ if $\text{Pr}(H=H_{1}|\bm{y}_{n})\geq c$ for a high value of $c\in(0,1).$ The higher $c$ is, the more confident is the decision. To see this, one simply observes that $(1-c)$ is the upper bound of posterior probability of a wrong rejection since when $H_{0}$ is rejected, $\text{Pr}(H=H_{0}|\bm{y}_{n})<1-c$.
@@ 3.2 Evidence
In simple words, evidence $e$ is the observed effect size from the trial data before they are observed. This means that in order to apply BESS, investigator needs to pre-specify (and calibrate) the potential observed effect size before the trial is conducted. This is analogous to the requirement of specifying the true parameter values in the standard SSE, except BESS assumes what might be observed rather than what might be true. We consider evidence $e$ as a function of the sufficient statistic in Table 1. For the three outcomes in one-arm trials, $(e+\theta_{0})$ is exactly the sufficient statistic. For two-arm trials, $e$ is the difference between the sufficient statistics $(\bar{y}_{0},\bar{y}_{1})$ of the two arms. Being a function of sufficient statistics allows for a search algorithm to find an appropriate sample size under BESS, which will be clear next. Another option is to define evidence as a test statistics, such as the $z$-statistics. However, this may reduce the interpretability of the evidence to nonstatisticians and may lead to more complicated calibration. We leave this option for future research.
@@ 3.3 Sample Size of BESS
For binary and count data in two-arm trials, $\text{Pr}(H=H_{1}|\bm{y}_{n})=\text{Pr}(H=H_{1}|\bar{y}_{1},\bar{y}_{0},n)$. See Appendix A.3 for a quick proof. Since $e=(\bar{y}_{1}-\bar{y}_{0})$, fixing evidence $e$ does not uniquely define $\text{Pr}(H=H_{1}|\bm{y}_{n}).$ However, for binary outcomes with $n$ samples per arm, $\bar{y}_{1}$ and $\bar{y}_{0}$ can only take on a finite number of values. Hence, for binary outcomes, we propose to specify evidence $e$ and find all pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$ that satisfies $e=(\bar{y}_{1}-\bar{y}_{0})$. We then find the minimum posterior probability of $H_{1}$ among these pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$, i.e.,
@@
Then the sample size statement for a two-arm trial is the same as BESS 1. And BESS Algorithm 2 provides details of finding the sample size $n$ for a two-arm trial for binary outcomes.
BESS Algorithm 2 Two-Arm Trials; Binary
@@
For count-data outcome, and as an alternative approach for binary outcomes, one may wish to specify the values of $\bar{y}_{1}$ and $\bar{y}_{0}$ directly instead of their difference $e$. Since $(\bar{y}_{1},\bar{y}_{0})$ is sufficient, the sample size search algorithm becomes easier. To this end, we propose a simple variation BESS Algorithm 2’ in Appendix A.4 and statement BESS 2.2 assuming $(\bar{y}_{1},\bar{y}_{0})$ are given.
@@
Lastly, since binary and count data are integer-valued, not all specified evidence value $e$ can be achieved for a given sample size in the proposed BESS Algorithms 1, 2 and 2’. For example, when the outcome is binary and sample size $n=10$, it is impossible to observe evidence $e=0.15$ since this requires having $n\cdot e=1.5$ more responders in the treatment arm than the control. The number of responders cannot be a fraction. To this end, we propose to round down the specified $e$ to the nearest possible value for a given $n$, i.e., $\frac{1}{n}\lfloor n\cdot e\rfloor,$ where $\lfloor$.$\rfloor$ . ⌋ is the floor function. This gives a conservative estimate of the sample size since even with smaller evidence, the sample size would still ensure the needed confidence.